i need to write an equation of the line containing the given point and parallel to given line i do not understand how to express in y=mx+b.
(2,9,); x+4y=9

Answer by dabanfield(803)

i need to write an equation of the line containing the given point and parallel to given line i do not understand how to express in y=mx+b.
(2,9,); x+4y=9
We need to rewrite the equation x + 4y = 9 in slope-intercept form:
x + 4y = 9
4y = -x - 9
y = (-1/4)*x - 9/4.
So from this we know that the slope of the line is -1/4. A line parallel to this line will also have a slope of -1/4.
So the equation of the parallel line is:
y = (-1/4)*x + b.
We need to find b. Since the point (2,9) is on our line we know:
9 = (-1/4)*2 + b
b = 9.5
The equation of the parallel line then is y = (-1/4)*x + 9.5